Sound travels at approximately 1500m/s in water
How long will it take a whale song to travel 100km?(dont forget to change the distance into meters.)
1
Expert's answer
2019-01-15T09:02:11-0500

The distance traveled by sound

l=vt

So, time

t=l/v=(100 000 m)/(1 500 m/s)=66.7 s